# Okapi: Efficiently Safeguarding Speculative Data Accesses in Sandboxed Environments

### v2

@@ 5. Trust Domain5
    Okapi eliminates Meltdown-style vulnerabilities by design since it only permits access to physical memory addresses in case of a valid, i.e., non-excepting, virtual address translation.
  6. Trust Domain6
@@ 7. Trust Domain7
The majority of the described extensions require only minor changes to the pipeline, like adding single bits to micro-operations or TLB entries. Okapi, to a great extent, re-uses the existing logic in the LSQ and requires no modification to the cache and memory subsystem. Implementing the Okapi instructions and the logic to determine the visibility point are the only significant hardware extensions.
6-B Additional Software Support
@@
  • Transitions from one sandbox below the size of a thread to another must perform an OkapiReset.
@@
The trusted entity, e.g., the operating system, needs to provide page-alignment for the data of each sandbox. The software developer can take care of the above measures by annotating secret data and using an extended compiler to insert OkapiReset and/or OkapiLoad instructions accordingly. In case the user wishes to increase the sections in $p_{H}$, the absence of gadgets can be accomplished by manual inspection or gadget analysis tools, e. g. [46, 22].
7 Performance Evaluation
The proposed architecture is implemented and evaluated based on the out-of-order CPU model o3 CPU provided by the gem5 simulator [6], configured according to Tab. I. First, we evaluate the performance of Okapi after implementing only the hardware features. This determines the overhead for enforcing breakout security in thread-sized sandboxes and limiting the gadget space for poisoning attacks to code pages that are accessed by the sequential data flow (Sec. 7-A). Next, we showcase how different jobs within the same thread of a WebAssembly runtime can easily be hardened against sandbox breakout attacks using the OkapiReset instruction (Sec. 7-B). Afterward, we examine the performance overhead caused by using the OkapiReset instruction to explicitly vault sensitive data in a cryptographic library (Sec. 7-C).
TABLE I: gem5 configuration
@@ 7 Performance Evaluation
7-A Thread-Level Sandboxing
Trust Domain2222333344445555Normalized execution timeover insecure baseline designNaïve DelayEager DelayDIFTOkapiGeomeanperlbench0perlbench1gcc0gcc1gcc2bwaves0bwaves1mcfcactuBSSNlbmxalancbmkdeepsjengimagickleelanabexchange2fotonik3droms0.80.80.80.811111.21.21.21.21.41.41.41.41.61.61.61.6
Figure 5: SPEC CPU2017 performance results
We benchmark the gem5 Okapi design by running the SPEC CPU2017 [8] reference workloads in full system mode using SimPoints [25]. The first 100 billion instructions are profiled to create a maximum of five SimPoints comprising 100 million measurement and 50 million warm-up instructions each for all workloads. The evaluation considers only the workloads that we were able to run on the unmodified o3 CPU model.
For comparison, besides Okapi, we implemented also other architectures that mitigate TES attacks targeting data at rest:
  • Naïve delay [37] – executes loads once they reach the head of the ROB.
  • Eager delay [37] – executes load instructions as soon as they are no longer speculative.
  • A dynamic information flow tracking (DIFT) architecture similar to [48, 31] – In DIFT, speculative load instructions are allowed to access the memory and initiate a taint for the accessed data. Speculative instructions that are able to form side channels are blocked from using tainted data.
For the experiments reported in this subsection neither OkapiReset nor OkapiLoad instructions were inserted into the benchmarks. The performance results for the SPEC CPU2017 suite are depicted in Fig. 5, summarizing the runtime overhead normalized to the insecure baseline design. It can be seen that Okapi outperforms all other mitigation schemes with an average overhead of only 3.17 %. The easy-to-implement blanket fixes, naïve and eager delay, incur a delay of roughly 2.6$\times$ or 2$\times$, respectively. Although it needs more complex modifications in the processor, the DIFT approach still imposes an overhead of 18.36 %. The significant performance increase originates in Okapi’s nature of adding an entire data page to the trust domain at a time. This means that it does not need to delay loads to the same page twice. The other examined fixes enforce weak speculative non-interference on address level. However, all architectures provide the same level of breakout security if the thread’s data is page-aligned.
@@ 7-A Thread-Level Sandboxing
Figure 6: Okapi: dispatched load instructions
[…]
Figure 7: Okapi: first-time issued load instructions
In the following, we analyze the results for the Okapi architecture in more detail. The vast majority of the load instructions enters the ROB as potentially malicious, as depicted in Fig. 6. On average, 85.5 % of the dispatched load instructions commit their results to the architectural state. However, this varies from 60.7 % in the case of mcf to 99.9 % for cactuBSSN. Fig. 7 illustrates the status of speculative load instructions when the LSQ first tries to execute their address translation. A load instruction can either be blocked due to being marked as suspicious_load or be passed to the TLB. On average, around 14.3 % of the speculative load instructions are blocked from going to the TLB due to our countermeasure to prevent the use of Spectre-BTB gadgets. Out of the remaining load instructions that are passed to the TLB speculatively, 98.4 % have their safe access bit set, confirming our locality assumption.
@@
7-B Mutually Distrusting Sandboxes
We modified the Wasmtime WebAssembly runtime [10] to execute an OkapiReset instruction upon switching from one job to another. Resetting the permissions gained by the previously executed tenants effectively disables the possibility to mount a speculative breakout attack, even when multiple jobs are sharing the same thread.
@@ 7-B Mutually Distrusting Sandboxes
In this experiment we split the Shootout workloads from the Sightglass Benchmark Suite [9] into two sets and reduce the iterations performed by the more computation intensive workloads to keep the simulation effort feasible. For each set, we set up the Wasmtime environment to execute the workloads interleaved. We run the two sets on our Okapi CPU as well as on an unmodified baseline CPU for comparison.
On average, Okapi incurs an overhead of only 1.68 % to mitigate breakout attacks between the different tenants.
7-C Vaulting
TABLE II: Libsodium Results
Config atomic $p_{H}$ function-sized $p_{H}^{\prime}$
runtime overhead #reset runtime overhead #reset

### v3

@@ 6.1. Hardware Architecture
  6. (6)
@@
The majority of the described extensions require only minor changes to the pipeline, such as adding single bits to micro-operations or TLB entries. Okapi, to a large extent, re-uses the existing logic in the LSQ and requires no modification to the cache and memory subsystem. Implementing the Okapi instructions and the logic to determine the visibility point are the only significant hardware extensions.
6.2. Additional Software Support
@@
  • Transitions from one software sandbox below the size of a thread to another must perform an OkapiReset.
@@
The trusted entity, e. g., the operating system, needs to provide page alignment for the data of each sandbox. The software developer can implement the above measures by annotating secret data and using an extended compiler to insert OkapiReset and/or OkapiLoad instructions accordingly. In case the user wishes to increase the sections in $p_{H}$, the absence of gadgets can be verified by manual inspection or gadget analysis tools, e. g. (Wiebing et al., 2024; Google Security Research, 2020).
7. Performance Evaluation
The proposed architecture is implemented [^2] Available on GitHub: https://github.com/RPTU-EIS/gem5-Okapi and evaluated based on the out-of-order CPU model o3 CPU provided by the gem5 simulator (Binkert et al., 2011), configured according to Tab. 1. First, we evaluate the performance of Okapi after implementing only the hardware features. This determines the overhead for enforcing breakout security in thread-sized sandboxes and limiting the gadget space for poisoning attacks to code pages that are accessed by the sequential data flow (Sec. 7.1). Next, we showcase how different jobs within the same thread of a WebAssembly runtime can easily be hardened against sandbox breakout attacks using the OkapiReset instruction (Sec. 7.2.1). Afterwards, we showcase how the MPK-based sandboxing scheme ERIM (Vahldiek-Oberwagner et al., 2019) can be adapted to Okapi (Sec. 7.2.2). Lastly, we examine the performance overhead caused by using the OkapiReset and OkapiLoad instructions to explicitly vault sensitive data in a cryptographic library (Sec. 7.3).
Table 1. gem5 configuration
@@ 7. Performance Evaluation
7.1. Thread-Level Sandboxing
@@
This figure shows the SPEC CPU2017 performance results of Okapi and three countermeasures that are discussed in the literature. These countermeasures are Naïve Delay and Eager Delay described in (Sakalis et al., 2019), and dynamic information flow tracking (DIFT) similar to (Yu et al., 2019; Loughlin et al., 2021). The performance results are given as normalized execution time over the insecure baseline design. The results show that Okapi incurs about 3.17 % overhead in the geometric mean and outperforms all other approaches for nearly all benchmark programms. DIFT is the runner-up with 18.36 % overhead on average.
Figure 5. SPEC CPU2017 performance results
We run the SPEC CPU2017 (Bucek et al., 2018) reference workloads in full system mode using SimPoints (Hamerly et al., 2005) to benchmark the gem5 Okapi design. The first 100 billion instructions are profiled to create a maximum of five SimPoints comprising 100 million measurement and 50 million warm-up instructions for each workload. The evaluation considers only the workloads that we were able to run on the unmodified o3 CPU model.
For comparison, in addition to Okapi, we also implemented other hardware-based architectures that mitigate TES attacks targeting data at rest:
  • Naïve delay (Sakalis et al., 2019) – executes loads once they reach the head of the ROB.
  • Eager delay (Sakalis et al., 2019) – executes load instructions as soon as they are no longer speculative.
  • A dynamic information flow tracking (DIFT) architecture similar to (Yu et al., 2019; Loughlin et al., 2021)333 STT (Yu et al., 2019) reports numbers based on SPEC CPU2006. The provided code does not run on the latest versions of gem5. Therefore, we ported STT to gem5 23.0.0.1, so that all the compared schemes run the same checkpoints of the most recent SPEC CPU benchmarks on the same simulator. The performance overhead reported by the authors of STT using the same sources of speculation is 14.5 %. – In DIFT, speculative load instructions are allowed to access the memory and initiate a taint for the accessed data. Speculative instructions that are able to form side channels are blocked from using tainted data.
For the experiments reported in this subsection neither OkapiReset nor OkapiLoad instructions were inserted into the benchmarks. The performance results for the SPEC CPU2017 suite are depicted in Fig. 5, summarizing the runtime overhead normalized to the insecure baseline design. It can be seen that Okapi outperforms all other mitigation schemes with an average performance overhead of only 3.17 %. The easy-to-implement blanket fixes, naïve and eager delay, incur a delay of 162.03 % or 98.44 % , respectively. While DIFT requires more complex modifications to the processor’s hardware (cf. Sec. 11), it still imposes a performance overhead of 18.36 %.
The significant performance advantage of Okapi originates in its nature of adding an entire data page to the trust domain at a time. This means that it does not need to delay loads to the same page twice. The other examined fixes enforce weak speculative non-interference at address level. However, all architectures provide the same level of breakout security if the thread’s data is page-aligned.
In the following, we analyze the results for the Okapi architecture in more detail. The vast majority of the load instructions enters the ROB as potentially malicious, as depicted in Fig. 9 Appendix A. On average, 85.5 % of the dispatched load instructions commit their results to the architectural state. However, this varies from 60.7 % in the case of mcf to 99.9 % for cactuBSSN. Fig. 10 in Appendix A illustrates the status of speculative load instructions when the LSQ first tries to execute their address translation. A load instruction can either be blocked due to being marked as suspicious_load or be passed to the TLB. On average, around 14.3 % of the speculative load instructions are blocked from going to the TLB due to our countermeasure to prevent the use of Spectre-BTB gadgets. Out of the remaining load instructions 98.4 % have their safe access bit set, confirming our locality assumption.
@@
7.2. Mutually Distrusting Sandboxes
This subsection showcases how two state-of-the-art sandboxing frameworks can easily be extended with Okapi to enforce additional in-process isolation for the speculative domain.
7.2.1. WebAssembly
We modified the Wasmtime WebAssembly runtime (Bytecode Alliance, 2017) as shown in Fig. 6. In this use case, the trusted runtime executes a set of untrusted workloads (jobs) in a round robin schedule. Whenever a job runs out of so called fuel, it hands back the control to the runtime (black arrows). Every time the runtime calls or suspends a job, it executes an OkapiReset instruction (green arrows). This ensures that the newly assigned job has no speculative access rights. As a result, a malicious job can neither transiently access another job’s data nor the data belonging to the runtime (red arrows).
@@ 7.2.1. WebAssembly
In this experiment we use the Shootout workloads from the Sightglass Benchmark Suite (Bytecode Alliance, 2019) and reduce the iterations performed by the more computation-intensive workloads to keep the simulation effort feasible. We select the benchmarks with a CPU time greater than one second to capture the most realistic workloads with a representative amount of switches and OkapiReset instructions. We set up the Wasmtime environment to execute the workloads interleaved. We run the benchmarks with a varying amount of fuel on our Okapi CPU as well as on an unmodified baseline CPU for comparison.
The results are depicted in Fig. 7. On average, Okapi incurs an overhead ranging from 5.52 % for ten thousand fuel to only 2.34 % for one million fuel to mitigate breakout attacks between the different tenants. The experiment shows that running the same workloads with more fuel, i.e., fewer trust domain switches is beneficial for the Okapi architecture.
[…]
The figure visualizes the switch between two WebAssembly Runtime jobs. Executing the OkapiReset instruction whenever switching from the WebAssembly runtime to a job effectively blocks Spectre attack vectors between running jobs as well as Spectre attack vectors that target the WebAssembly runtime from one of the jobs.
Figure 6. WebAssembly runtime environment extended with OkapiReset instructions when starting/continuing execution of a job (green). This mitigates the Spectre attack possibilities (red) between the trust domains (cyan) and the runtime.
[…]
The figure shows the Shootout performance results of Okapi. The workloads are run interleaved with a varying amount of fuel. The results show that Okapi incurs about 2.34 % overhead in the geometric mean for one million fuel and 5.52 % for ten thousand fuel.
Figure 7. Sightglass Shootout
7.2.2. ERIM
In a second case study, we investigate the security synergies of Okapi with the ERIM (Vahldiek-Oberwagner et al., 2019) sandboxing library. Okapi prevents speculative execution of code located on different memory pages by its hardware and can therefore extend ERIM with execution prevention in case of speculatively misguided control flows, as, e.g., in a Spectre-BTB or Spectre-RSB attack.
Additionally, we instrument the erim_switch_to_untrusted macro with our newly introduced OkapiReset instruction, as depicted in Fig. 8. The erim_switch_to_trusted macro is adapted analogously. Note that no other modification of the framework is needed to extend ERIM to Okapi. While the memory protection key prevents the cross-domain access, resetting the safe access bits restricts the speculatively accessible memory space. This ensures that previously accessed memory regions accessible by the current set of memory keys cannot be leaked speculatively within an ERIM domain.
In our experiment we run a bubblesort algorithm for arrays of 64 and 4096 bytes. Adding OkapiReset to the ERIM domain switches does not affect the runtime of the algorithm.
1 #define erim_switch_to_untrusted
2do {
3 ERIM_SWITCH_TO_UNTRUSTED_STACK;
4 __wrpkrucheck(ERIM_UNTRUSTED_PKRU);
5 ERIM_INCR_CNT(1);
6 asm volatile("OkapiReset");
7} while(0)
[…]
The listing shows the code that is being executed when ERIM switches to the untrusted domain. Adding Okapis protection guarantees is as simple as adding the OkapiReset instruction at the end of the code.
Figure 8. Okapi enhanced ERIM switch to the untrusted domain
7.3. Vaulting
The previous subsections consider security against breakout attacks. They show that Okapi (i) performs well for a representative set of workloads (cf. 7.1), (ii) introduces only minor performance overhead when applied in realistic in-process isolation scenarios (cf. 7.2.1), and (iii) requires only minor software modifications when combined with memory protection keys for in-process isolation (cf. 7.2.2).
This subsection addresses the remaining case of poisoning attacks. We benchmark the Libsodium (Denis, 2023) constant-time cryptographic library which already comes with a page-aligned memory allocator. This feature facilitates achieving poisoning security (cf. Sec. 6.2).
In this experiment, we compare three different implementations of security-sensitive code sections $p_{H}$ which we refer to as atomic $p_{H}$, OkapiLoad $p_{H}$, and function-sized $p_{H}^{\prime}$. For atomic $p_{H}$ we instrument the hashing, secret and public key routines within the cryptographic library to execute an OkapiReset instruction after every load instruction targeting secret data to prevent poisoning attacks from leaking secret data. OkapiLoad $p_{H}$ achieves the same level of security by replacing every load instruction to secret data with an OkapiLoad instruction. This instrumentation allows all sections in $p_{L}$ to speculate freely on data accessed by their predecessors. Lastly, function-sized $p_{H}^{\prime}$ adapts the same routines to reset the permissions on function returns. This instrumentation approximates the security-critical parts of the program to the set of sections $p_{H}^{\prime}$ consisting of the code page(s) of entire functions. While this measure may not guarantee full security, it drastically reduces the attack surface as well as the potential gadget space in proportion to the region size (cf. Sec. 9.2). This can be automated w. r. t. to a predefined granularity. To this end, we modified the LLVM compiler framework (LLVM Foundation, 2007). LLVM’s intermediate representation allows for an easy adaptation of the chosen granularity.
Tab. 2 in Appendix A summarizes the number of Okapi instructions executed by the different instrumentations.
We performed 100 runs of different routines in Libsodium and normalized the median runtime to the unmodified version on an unmodified CPU. The finest possible granularity of $p_{H}$ implemented by OkapiReset incurs an average performance overhead of 61.72 % to achieve weak speculative non-interference. While maintaining the same level of security this overhead can be reduced to 59.78 % by using OkapiLoad. This is a clear improvement over software-based mitigations (Corporation, 2018a, c, b; Turner, 2018) currently rolled out in industry.
Resorting to the more coarse-grained function-sized $p_{H}^{\prime}$ significantly reduces the incurred overhead even further to 11.60 %.
For more detail on the individual overheads refer to Fig. 11 in Appendix A.
Note that the Okapi security features against poisoning complement the security level provided by the constant-time library. Without Okapi, a poisoning attack can redirect the control flow such that the vault leaks the secret despite constant-time measures being in place (cf. (Yavarzadeh et al., 2024)).
8. Hardware Overhead
We implemented the Okapi architecture according to Sec. 6 in the open-source out-of-order processor BOOM (Zhao et al., 2020). Afterwards, we performed synthesis and place & route in Synopsys DesignCompiler and IC Compiler II for both the original design and our OkapiBOOM (both in medium configuration) at 500 MHz on a commercial 12 nm FinFET process from GlobalFoundries. Okapi does not impact the critical path of the chip  allowing us to achieve the same maximum clock frequency as the original BOOM core. Compared to the baseline design, Okapi increases cell area by only 0.6 % while reducing the power consumption by 1.53 %. For additional details, refer to Appendix C. This confirms our claim that the Okapi architecture incurs a negligible hardware overhead, in particular, when compared to DIFT (cf. Sec. 11).
9. Security Evaluation
9.1. Formal Hardware Verification
In Appendix B we apply a standard Spectre attack to the gem5 o3 CPU with and without Okapi protection. Mounting a single example attack and showing it to be blocked by Okapi demonstrates the Okapi mechanisms but does not provide full confidence in the claimed security guarantees. Therefore, we also conducted a formal analysis to provide an exhaustive security proof for the Okapi architectural concept.
We base our verification on the methodology of Unique Program Execution Checking (UPEC), as presented in (Fadiheh et al., 2023), which formally establishes weak speculative non-interference for a design under verification (DUV). Our DUV is BOOM extended by the described Okapi architecture. [^4] Available on GitHub: https://github.com/RPTU-EIS/OkapiBOOM We adapt UPEC to sandboxing (cf. Sec. 3) by constraining the UPEC property (Fadiheh et al., 2023) such that the program in its sequential semantics never accesses high memory locations.
The Okapi-adapted UPEC property holds after a proof runtime of 41 minutes. This guarantees breakout security for sandboxes of custom size, under the assumption that the software executes an OkapiReset instruction on sandbox exits and that it separates confidential and public data pages.These assumptions can be easily checked for a given software. As the original BOOM design has been proven susceptible to Meltdown (Jauch et al., 2023), our formal verification also showcases the inherent mitigation of Meltdown by Okapi.
In addition, our formally verified hardware guarantees that any program $p$ is poisoning-secure if the software assumptions of Sec. 6.2 are fulfilled. This can be proven for a given software by checking that the program sections $p_{H}$ do not add the high memory locations to the trust domain.
9.2. Attack Surface Analysis
In this section, we discuss the reduction of the attack surface for Okapi in its least-overhead version, i.e., when only implementing its HW measures. While this version provides formally proven breakout security at thread level, the reduction of the attack surface for poisoning cannot be assessed formally. However, a quantitative analysis provides valuable insights.
As explained in Sec. 4.1, the number of speculatively accessible pages of a trust domain increases over time. However, the empirical results for SPEC CPU2017 shown in Fig. 2 demonstrate that programs access only 5.69 % of their mapped code pages and 31.27 % of their mapped data pages between two context switches [^5] Every context switch involves a privilege switch, on average. It is only this set of data pages visited sequentially between context switches that can be accessed speculatively by any Spectre gadget in the mapped address space. As Okapi without any software measures prevents the execution of gadgets following transient jumps to arbitrary code pages, this leads to a drastic reduction of the attack surface for poisoning since attackers could only exploit gadgets on the 5.69 % of the code pages that are reached by the program flow non-speculatively.
If developers of security-critical applications wish to remove also this last gap, they only need to care about Spectre gadgets in the immediate vicinity of locations (i. e., the pages) where the code works with secret data. Spectre gadgets in other locations are automatically suppressed by Okapi without any further action needed by the developers or the compiler.
An interesting observation we make is that, even if an attacker could successfully access confidential data, the cache side channel can only be exploited if the address used to encode the data also belongs to the trust domain. Hence, even without any software support, the Okapi hardware alone (besides guaranteeing breakout security) drastically increases the effort necessary for mounting a successful poisoning attack.
In conclusion, we believe that the hardware-only version of Okapi can be an attractive security solution in many practical settings. This and other sweet spots in Okapi’s trade-offs are summarized in the following section.
10. Sweet Spots in Okapi’s trade-offs
As presented in previous sections, Okapi allows for several trade-offs between the level of protection and associated cost. The range and characteristics of these trade-offs, unfortunately, cannot be easily formalized and simply mapped to formal notions of security. Instead, however, we have demonstrated in our experiments that different Okapi trade-offs match particularly well with certain practical settings. In the following, we point to such ”sweet spots” where Okapi, in practice, can help to achieve a high level of protection at exceptionally low cost:
  • Okapi only introduces an area overhead of 0.6 % after synthesis and place & route. For use cases without security requirements, the Okapi features can be switched off to avoid all performance overhead.
  • By its hardware measures alone, Okapi provides secure speculation for the sandboxing threat model, as formally defined in  (Guarnieri et al., 2021), at thread level.
    In addition, the OkapiReset instruction can be used to secure software sandboxes at a finer granularity. Our experiments in Sec. 7.2.1 show a performance overhead of only 1.68 % to eliminate the possibility for an untrusted Wasmtime workload to launch a breakout attack, even on workloads sharing the same thread.
  • As explained in the previous section, the above low-overhead measures do not only protect against sandboxing but, on top, provide significantly stronger security also for poisoning attacks. In addition, for systems without explicit sandboxing in software and for legacy code, this measure eliminates universal read gadgets and, more importantly, protects the kernel from being misused as a confused deputy.
  • There will remain special (arguably rare) situations where the remaining gadget space within one code page is considered insufficiently secure and the user wishes to achieve address-level weak speculative non-interference (Guarnieri et al., 2021). In this case, the OkapiLoad instruction fully closes this gap. The expected performance overhead may grow up to 60.0 %, which is fully competitive with current industrial solutions offering this level of security. However, such a high level of security is normally not necessary for the full application. Okapi allows for limiting its expensive measures to only those code sections where they are really needed, thus, in practice, avoiding excessive overhead over the full software.
  • Okapi enforces (in-)process isolation for the speculative domain and thereby complements designs that already implement other dedicated isolation mechanisms. Intel MPK (Corporation, 2024), for example, guards processes with additional memory protection keys, but the same keys are available to all threads within a process. This gap is avoided in Okapi by additionally providing speculative separation between threads. Other work (Schrammel et al., 2020) enables memory protection keys to also implement in-process isolation, but lack execution prevention. An attacker may still transiently divert the control flow to create a leakage. Such poisoning attacks are mitigated in Okapi by the mechanisms described in Sec. 4.2.
11. Related Work
SafeBet (Green et al., 2023) exploits temporal locality based on similar observations as in Okapi. However, the SafeBet mitigation scheme introduces a cache-like hardware structure to keep track of which code regions have already accessed which memory regions legally. The performance overhead incurred by SafeBet depends on the chosen granularity of the monitored code regions, and ranges from 6 % (at 1 GB granularity) to over 270 % (at address level). Also, the hardware overhead is significantly higher than that of Okapi, due to the additional 8.3 KB high-speed cache needed for access mapping.
SpecBox (Tang et al., 2023) mitigates cache-based TES attacks by hiding side effects of speculative loads within the cache at cache-line granularity with minimal performance overhead. Okapi, in comparison, protects against all TES attacks (including other, non-cache-based leakage channels, e.g., Spectre-STC (Fadiheh et al., 2020), and Speculative Interference Attacks (Behnia et al., 2021)) by already preventing the speculative access to protected data. Okapi protects data at a coarser granularity (page-level) which is well suited for sandboxing scenarios, such as FaaS, and therefore comes with less implementation overhead.
Hardware-only mitigations based on DIFT (Yu et al., 2020, 2019; Jauch et al., 2023; Choudhary et al., 2021; Loughlin et al., 2021) protect data at rest with a performance overhead ranging from 10.5 % (Yu et al., 2020) to 36.0 % (Jauch et al., 2023). Additional hardware protection for data in transit incurs a performance overhead of up to 45.0 % (Choudhary et al., 2021). While these techniques deliver security guarantees at address-level granularity, they come at the price of higher hardware and performance overheads compared to Okapi. Most importantly, they significantly increase the design complexity of the microarchitecture, particularly across the memory hierarchy, which impedes their adoption.
A second class of hardware-based mitigations utilizes manual annotations by the programmer to inform the hardware about what data to protect. These approaches range from mitigations for Spectre-PHT (Fustos et al., 2019) to schemes that protect all annotated data in transit against speculative leakage (Schwarz et al., 2020; Daniel et al., 2023). While SpectreGuard (Fustos et al., 2019) offers less protection than Okapi, ConTeXT (Schwarz et al., 2020) inherits the hardware overhead of DIFT approaches and incurs a performance overhead of up to 71.14 % in security-critical applications. ProSpeCT (Daniel et al., 2023) only applied a simplified version of its architecture to custom benchmarks and achieved a performance overhead of 10.0 % to 40.0 %.
Perspective (Kim et al., 2024) mitigates active as well as passive attacks on the OS kernel. Active attacks in Perspective’s taxonomy can be compared to Okapi’s breakout attacks while the passive attacks can be mapped to poisoning attacks. Perspective achieves similar security as Okapi but its scope is limited to the kernel while Okapi can mitigate attacks on and by smaller sandboxes even if they share the same process. Additionally, Perspective requires modifications to the kernel and a static or dynamic analysis of the system calls.
Okapi effectively enforces dynamic sandboxing for speculative data accesses and therefore shares this objective with conventional sandboxing approaches and (in-)process isolation techniques that also mitigate certain Spectre attacks (Narayan et al., 2021; Schrammel et al., 2020; Narayan et al., 2023; Bourgeat et al., 2019). However, these mitigations pursue a different security target and introduce a much higher overhead in implementation and performance. In addition, while most of these techniques can thwart Spectre-PHT attacks, many of them leave an open attack surface to Spectre-BTB.
Retpolines (Turner, 2018) are a software construct to prevent branch target injection. They pursue a similar goal as Okapi’s hardware mechanism to restrict Spectre attacks after speculative instruction page crossings. Since the initial introduction there has been effort to improve on the high performance overhead introduced by retpolines. In (Duta et al., 2021) the authors reduce the initial overhead from 149 % to 10.6 %. Retpolines require recompilation of the software that needs to be protected and provides incomplete protection on some CPUs (Wikner and Razavi, 2022).
Further, Intel rolled out (enhanced) Indirect Branch Restricted Speculation (IBRS) (Corporation, 2018b), Single Thread Indirect Branch Predictors (STIBP) (Corporation, 2018c) as well as Indirect Branch Predictor Barriers (IBPB) (Corporation, 2018a) in response to Spectre-BTB. Both, IBRS and STIBP have been shown to still be vulnerable to branch target injection attacks (Li et al., 2024). While IBPB is the most strict countermeasure, it introduces a high performance overhead when activated. Hence, it is not well suited for Function as a Service (FaaS) scenarios that require frequent security domain transitions (Li et al., 2024).
In 2019, Intel introduced a control-flow enforcement technology (CET) (Shanbhogue et al., 2019) that restricts the possible control flow after an indirect branch to legal jump targets. While this protects software from a control flow redirection to arbitrary targets, the set of possible targets still leaves space for poisoning attacks.
Serberus (Mosier et al., 2024) presents a methodology to harden constant-time code against Spectre attacks in software. The authors report overheads ranging from 0.9 % to 196.2 % with a standard deviation of around 80 % on a selection of five cryptographic primitives. It depends on existing control flow enforcement technologies and LLVM compiler features to provide a software-based protection of static constant-time code for the constant-time threat model (Guarnieri et al., 2021). While the results are promising we consider Serberus as an orthogonal approach.
12. Conclusion
At its core, Okapi prevents these attacks by delaying speculative accesses to data pages which have not yet been accessed from within the currently running sandbox. Okapi provides the possibility to customize the granularity of sandboxes, thereby covering a range of possible applications. The Okapi hardware architecture is easy to implement and can be seamlessly integrated into existing designs. In particular, breakout security, being highly relevant in applications of edge computing, is achieved with an exceptionally low hardware and performance overhead. This bears promise for Okapi’s adoption in industrial practice, possibly replacing costly software patches.
An interesting direction for further research could be the combination of the Okapi hardware mechanisms with orthogonal software protection mechanisms, e. g., Serberus (Mosier et al., 2024) for further performance improvements.
Acknowledgements.
We are grateful to all our reviewers and the shepherd for substantial feedback and help with revising the paper. We further thank Jan Lappas and Norbert Wehn for their valuable support. The reported research was partly supported by the Sponsor DFG SPP Nano Security https://www.dfg.de/en/news/news-topics/announcements-proposals/2023/info-wissenschaft-23-11 under grant Grant #HaSPro, by the Sponsor Agentur für Innovation in Cybersicherheit GmbH https://www.cyberagentur.de/ under grant Grant #PROTECT (EVIT program), by the Sponsor BSI https://www.bsi.bund.de under grant Grant #SILKOSTU and by the Sponsor Intel Corp. https://www.intel.de/ Grant #Scalable Assurance Program.
Appendix A Additional experimental Results
Fig. 9 shows the number of safe and unsafe loads that are dispatched during the SPEC CPU2017 benchmarks. It is visible that the vast majority of loads entering the ROB are unsafe loads of which on average 85.5 % are committed eventually.
Fig. 10 shows the status of speculative load instructions when the LSQ tries to execute their address translation for the first time. On average, around 14.3 % of the speculative load instructions are blocked from going to the TLB due to our countermeasure to prevent the use of Spectre-BTB gadgets. Out of the remaining load instructions that are passed to the TLB speculatively, 98.4 % have their safe access bit set.
Tab. 2 provides the exact number of Okapi instructions added by the different instrumentations. atomic $p_{H}$ adds an OkapiReset after every load instruction, while OkapiLoad $p_{H}$ replaces every load instruction by an OkapiLoad. Thus, both instrumentations have the same amount of special instructions. However, atomic $p_{H}$ has a larger binary size because it adds instructions instead of replacing them. The third option, function-sized $p_{H}^{\prime}$, has the least amount of added instructions.
Fig. 11 summarizes the runtime overhead of the median execution times of 100 runs of the different routines available in the Libsodium library normalized to the unmodified version on an unmodified CPU. As can be seen, the finest possible granularity of $p_{H}$ implemented by OkapiReset introduces an average performance overhead of 168.43 %. While maintaining the same level of security this overhead can be reduced to 59.78 % by using OkapiLoad. Resorting to the more coarse-grained function-granular choice of $p_{H}^{\prime}$ significantly reduces the incurred overhead even further down to 11.60 %
[…]
The figure shows the number of safe and unsafe loads that are dispatched during the SPEC CPU2017 benchmarks. It is visible that the vast majority of loads entering the ROB are unsafe loads of which on average 85.5 % are committed eventually.
Figure 9. Okapi: dispatched load instructions
[…]
The figure shows the status of speculative load instructions when the LSQ first tries to execute their address translation. On average, around 14.3 % of the speculative load instructions are blocked from going to the TLB due to our countermeasure to prevent the use of Spectre-BTB gadgets. Out of the remaining load instructions that are passed to the TLB speculatively, 98.4 % have their safe access bit set.
Figure 10. Okapi: first-time issued load instructions
Table 2. Libsodium Okapi Instructions
Config atomic $p_{H}$ OkapiLoad $p_{H}$ function-sized $p_{H}^{\prime}$
#OkapiReset #OkapiLoad #OkapiReset